OpenStudy (daysiphoto):

The probability that a student is accepted to a prestigious college is 0.3. If 5 students from the same school apply. What is the probability of the following. Use the Binomial Probability formula a. That at most 2 are accepted? b. That 2 or more but less than 4? c. That all 5 are accepted?

OpenStudy (harman.singh):

For a), you would need to use the binomial formula to find the probability for '1' and '2', and then add them together

OpenStudy (harman.singh):

Just to be certain, are you familiar with the Binomial Probability formula or used it before?

OpenStudy (daysiphoto):

have not used it... but do I plug in?

OpenStudy (daysiphoto):

http://www.mathwords.com/b/binomial_probability_formula.htm this is how I would do it correct?

OpenStudy (harman.singh):

Found this on the web. Might be useful for this question 

OpenStudy (harman.singh):

- p is the probability - x is the number of successful trials - n is the total number of events possible

OpenStudy (daysiphoto):

ok awesome thank you so much... let me see if I can work it out

OpenStudy (harman.singh):

Oh, I didnt see your early reply. But yes, the formula in that link is correct

OpenStudy (harman.singh):

Just to correct myself, for a), you would need to find the probability for '0', '1' and '2' . I forgot to add the zero before. My apologies.

OpenStudy (daysiphoto):

so I would have to use 0.3 and 5 as my numbers for the formula?

OpenStudy (harman.singh):

Yes, correct. So here is how it would look like: \[\left(\begin{matrix}n \\ k\end{matrix}\right)*p^k * (1-p)^(n-k)\]

OpenStudy (harman.singh):

So when k=1, \[\left(\begin{matrix}5 \\ 1\end{matrix}\right)*0.3^1*(1-0.3)^(5-1)\]

OpenStudy (daysiphoto):

ok working on...

OpenStudy (daysiphoto):

= 4.2 ?

OpenStudy (harman.singh):

Umm..no. I am not exactly sure where you went wrong. This is what you should get for each one for a): p(0) = 0.16807 p(1) = 0.36015 p(2) = 0.3087 After adding these, I get a final answer of 0.83692. Are you sure you used the formula correctly?

OpenStudy (harman.singh):

Here is how I found 'p(1)' step-by-step using the formula So we are given that, k=1 p=0.3 n=5 \[\left( \frac{ 5! }{ 1!(5-1)! } \right)*0.3^1*(1-0.3)^{5-1}\] \[\left( \frac{ 120 }{ 24 } \right)*0.3*0.7^{4}\] This gives me an answer of 0.16807 for 'p(1)'. Hopefully, this can help you figure out where you made the mistake in your calculations. Let me know if you need more explanation.

OpenStudy (daysiphoto):

I see what I did wrong how did you solve it?

OpenStudy (harman.singh):

Solve which part?

OpenStudy (daysiphoto):

I see how you did it, sorry I am so lost but I am going to write it out again

OpenStudy (harman.singh):

Thats fine

OpenStudy (daysiphoto):

b(x < 2; 5, 0.3) = b(x = 0; 5, 0.3) + b(x = 1; 5, 0.3) + b(x = 2; 5, 0.3) b(x < 2; 5, 0.3) = 0.1681 + 0.3601 + 0.3087 b(x < 2; 5, 0.3) = 0.8369

OpenStudy (harman.singh):

Awesome! Looks correct. Here is more information to solve other parts of the question: For b), it says 2 or more but less than 4. That means its asking for probability when k=2 & 3. For c), all 5 are accepted meaning find the probability when k=0,1,2,3,4, & 5. Good Luck!

OpenStudy (daysiphoto):

Thanks I got it!

OpenStudy (harman.singh):

My pleasure! Glad I could help.